OpenStudy (anonymous):

In solving the equation (x + 3)(x – 3) = -8, Eric stated that the solution would be x + 3 = -8 => x = -11 or (x – 3) = -8 => x = -5 However, at least one of these solutions fails to work when substituted back into the original equation. Why is that? Please help Eric to understand better; solve the problem yourself, and explain your reasoning.

OpenStudy (anonymous):

x^2 - 9 = 8

OpenStudy (anonymous):

Is that what you're trying to do?

OpenStudy (anonymous):

I don't understand it and I have to show step by step

OpenStudy (anonymous):

Okay, are you trying to solve for x?

OpenStudy (anonymous):

I believe it is trying to find out which one is incorrect at the soluitions and why it fails

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

I don't see anywhere in the equation that talks about inequalities so I'll do solve it my way. Is that okay?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

The reason Eric was getting the wrong answer is because - 5 cannot be an answer. Even though - 11 is an answer, you cannot take two variables that are being multiplied and set them to equal zero without distributing first. You must first work out the equation so that x can be manipulated to get a constant value. For instance, (x-3)(x+3) =8 Subtract the 8 to the left side: x^2 - 17 = 0 now add 17 to the right side to get x by itself. Because x is squared, you must take the square root of both sides. X will the be equal to : \[- and + \sqrt{17}\]

OpenStudy (anonymous):

Does that make any sense?

OpenStudy (anonymous):

yes thank u so much

OpenStudy (anonymous):

No problem :D